# Robust $Q$-Learning for finite ambiguity sets

### v2

@@ 6.1. Auxiliary Results
For any function $f:\mathcal{X}\times A\rightarrow\mathbb{R}$, we write
(6.1) $\|f\|_{\infty}:=\sup_{x\in\mathcal{X}}\sup_{a\in A}|f(x,a)|.$
Next, we introduce the operator $\mathcal{H}$ operating on a function $v:\mathcal{X}\times A\rightarrow\mathbb{R}$ being defined by
(6.2) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}v(x,a):=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}v(X_{1},b)\right].$
@@
Let $10<\alpha<$ 1, and let the ambiguity set $\mathcal{P}$ be defined in (2.1). Then the following fixed-point equation holds true for the optimal $Q$-value function defined in (2.5)
@@
This follows directly by definition of $Q^{*}$ and by Proposition 2.1, compare also [29, Proof of Lemma 17]. ∎
@@
The proof follows is analogue to [29, Proof of Lemma 18]. ∎
The following lemma is a result from stochastic approximation theory adjusted to our setting, compare [29, Lemma A.1] for the presented formulation adjusted to the underlying setting. We refer also to [39, Lemma 1], [15] [19, Theorem 1], [41, Lemma 12], and [46, Lemma 3] for further reference.
@@
    There exists $\delta\in(0,1)$ such that $\|\mathbb{E}_{\mathbb{P}_{0}}\left[F_{t}(\cdot,\ \cdot)~{}\middle|~{}\mathcal{G}_{t}\right]\|_{\infty}\leq\delta\|\Delta_{t}\|_{\infty}$ $\mathbb{P}_{0}$-almost surely for all $t\in\mathbb{N}_{0}$.
@@
Let $Z$ be a random variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ satisfying $m\leq Z\leq M$ for some $-\infty<m\leq M<\infty$. Then, we have
@@ 6.2. Proofs of the results from Section 2, Section 3, and Section 4
Note that the equality $\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)$ directly follows by definition of the operator $\mathcal{T}$, and it only remains to show $\mathcal{T}V=V$. To this end, we aim at applying [31, Theorem 3.1] with $p=0$ and verify its assumptions, i.e., we verify [31, Assumption 2.2] and [31, Assumption 2.4] by showing the following four properties.
@@
    $(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a):=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\}$
@@
To see that (i) holds, we note that $\mathcal{P}$ is by definition non-empty and since the set $\mathcal{P}(x,a)$ is finite it is also compact for all $(x,a)\in\mathcal{X}\times A$. To show the upper hemicontinuity, we consider some $(x,a)\in\mathcal{X}\times A$ and we let $(x_{n},a_{n},\mathbb{P}_{n})_{n\in\mathbb{N}}\subseteq\operatorname{Gr}(\mathcal{P}):=\{(x,a,\mathbb{P})~{}|~{}(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)\}$ be a sequence such that $(x_{n},a_{n})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$. Since both $\mathcal{X}$ and $A$ are finite, we have $(x_{n},a_{n})=(x,a)$ for $n$ large enough. Hence, $\mathbb{P}_{n}\in\mathcal{P}(x,a)$ for $n$ large enough. As $\mathcal{P}(x,a)$ is finite we also have that $\mathbb{P}_{n}=\mathbb{P}^{(i)}(x,a)$ for infinitely many indices $n$ for some $i\in\{1,\dots,N\}$. Hence, we can find a subsequence $(\mathbb{P}_{n_{k}})_{k\in\mathbb{N}}$ such that $\mathbb{P}_{n_{k}}\rightarrow\mathbb{P}^{(i)}(x,a)\in\mathcal{P}(x,a)$ weakly as $k\rightarrow\infty$ and the upper hemicontinuity of $\mathcal{P}$ follows by the characterization provided in [2, Theorem 17.20].
@@
and we obtain, since $(x_{n},a_{n})=(x,a)$ for $n$ large enough that $\mathbb{P}_{n}\rightarrow\mathbb{P}$ weakly which shows the lower hemicontinuity of $\mathcal{P}$ according to [2, Theorem 17.21].
@@
Finally, (iv) follows since $C_{P}=1$ and as we have $10<\alpha<$ 1 by assumption. ∎
@@
To establish convergence of the $Q$-learning algorithm from Algorithm 1, we aim at applying Lemma 6.3. To verify the assumptions of Lemma 6.3 we first note that for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ the update rule from (3.2) can be written equivalently as
@@
(6.5) $\gamma_{t}(x,a,X_{t}):=1\mathrm{l}_{\left\{(x,a)=\left(X_{t},~{}a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{t}.$
Next, for for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ we define
$\Delta_{t}(x,a):$ $=Q_{t}(x,a)-Q^{*}(x,a),$
$F_{t}(x,a):$ $=\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right)1\mathrm{l}_{\left\{(x,a)=\left(X_{t},~{}a_{t}(X_{t})\right)\right\}}.$
@@
$k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,...,N}\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right].$
Once we have verified that the three assumptions of Lemma 6.3 are fulfilled, we obtain $\Delta_{t}(x,a)\rightarrow 0$ $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely as $t\rightarrow\infty$ and hence
$Q_{t}(x,a)\rightarrow Q(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$
@@
$\Delta_{t+1}(x,a)=$ $Q_{t}(x,a)-Q^{*}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big{(}r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)\Big{)}$
@@
$+\gamma_{t}(x,a,X_{t})\cdot\Big{(}r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)+Q^{*}(x,a)-Q^{*}(x,a)\Big{)}$
@@
$+\gamma_{t}(x,a,X_{t})\cdot\Big{(}r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\Big{)}$
@@
where we used in the last step that the indicator $1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}$ appears both in the definition of $F_{t}$ and $\gamma_{t}$ and that the square of the indicator function is the indicator function itself. Consider the filtration $(\mathcal{G}_{t})_{t\in\mathbb{N}}$ with
$\mathcal{G}_{t}:=\sigma({X_{1},X_{2},...,X_{t}}),\ t\in\mathbb{N},$
and $\mathcal{G}_{0}:=\{\varnothing,\Omega^{N}\}$ being the trivial sigma-algebra. Then, by definition we have for all $t\in\mathbb{N}$ and for all $(x,a)\in\mathcal{X}\times A$ that the random variables $\Delta_{t}(x,a),\ \gamma_{t}(x,a),\ F_{t-1}(x,a)$ are $\mathcal{G}_{t}$-measurable. Next, let $(x,a)\in\mathcal{X}\times A,$ and $t\in\mathbb{N}$. Then, we have by the definition of $\mathbb{P}_{x_{0},\textbf{a}}$ in (3.3) $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely that

### v3

@@ 7.1. Auxiliary results
We start this section by introducing necessary notation. For any function $f:\mathcal{X}\times A\rightarrow\mathbb{R}$, we write
(7.1) $\|f\|_{\infty}:=\sup_{x\in\mathcal{X}}\sup_{a\in A}|f(x,a)|,$
and we define an estimation of the ambiguity set after $t$ steps of Algorithm 1 via the set-valued map
(7.2) $\mathcal{X}\times A\ni(x,a)$ $\twoheadrightarrow\mathcal{P}_{t}(x,a)=\left\{\mathbb{P}_{t}^{(1)}(x,a),\cdots,\mathbb{P}_{t}^{(N)}(x,a)\right\}\subset\mathcal{M}_{1}(\mathcal{X}).$
where $\mathbb{P}_{t}^{(k)}$ is defined in (3.2). Next, we introduce the operators $\mathcal{H}_{t}$ and $\mathcal{H}$ operating on a function $q:\mathcal{X}\times A\rightarrow\mathbb{R}$ being defined by
(7.3) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}_{t}q(x,a):=\min_{\mathbb{P}\in\mathcal{P}_{t}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right]\in\mathbb{R},$
as well as
(7.4) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}q(x,a):=\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right]\in\mathbb{R}.$
Moreover, analogue to (2.5), we define the optimal $Q$-value function with respect to the ambiguity set $\mathcal{P}_{t}$ via
(7.5) $\mathcal{X}\times A\ni(x,a)\mapsto Q_{t}^{*}(x,a):=\min_{\mathbb{P}\in\mathcal{P}_{t}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
where - by slight abuse of notation - $V$ is defined as in (2.3) with using the ambiguity set $\mathcal{P}_{t}$.
@@
Let $0<\alpha<1$, and let the ambiguity set $\mathcal{P}_{t}$ be defined in (7.2). Then the following fixed-point equations hold true for the optimal $Q$-value functions defined in (2.5) and (7.5), respectively
$\mathcal{H}_{t}Q_{t}^{*}(x,a)$ $=Q_{t}^{*}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A,\text{ for all }t\in\mathbb{N},$
@@
This follows directly by definition of $Q_{t}^{*}$ and by Proposition 2.1, compare also [39, Proof of Lemma 17]. ∎
@@
$\|\mathcal{H}_{t}q_{1}-\mathcal{H}_{t}q_{2}\|_{\infty}$ $\leq\alpha\|q_{1}-q_{2}\|_{\infty}\text{ for all }t\in\mathbb{N},$
@@
The proof follows analogue to [39, Proof of Lemma 18]. ∎
Next, recall the definition
$\mathbb{P}_{t}^{(k)}(x,a)(y)=\frac{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}),X_{s+1}^{(k)})=(x,a,y)\}}}{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}}.$
Lemma 7.3.
Let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
(7.6) $\sum^{\infty}_{t=0}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}=\infty,\text{ for all }(x,a)\in\mathcal{X}\times A,\ \mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
Then, we have that
$\mathbb{P}_{t}^{(k)}(x,a)\rightarrow\mathbb{P}^{(k)}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A,k\in\{1,\dots,N\}\qquad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
as $t\rightarrow\infty$.
Proof.
Let $k\in\{1,\dots,N\}$ and let $(x,a)\in\mathcal{X}\times A$. First note that the condition in (7.6) ensures that
(7.7) $N_{t}(x,a):=\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}\rightarrow\infty~~\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely as }t\rightarrow\infty.$
Let $\tau_{j}(x,a)$ be the time of the $j$-th visit of $\left((X_{t},a_{t}(X_{t})\right)_{t\in\mathbb{N}}$ to $(x,a)$. Then, we draw $Y_{j}^{(k)}:=X_{\tau_{j}(x,a)+1}^{(k)}\sim\mathbb{P}^{(k)}(x,a)$. By construction $(Y_{j}^{(k)})_{j\in\mathbb{N}}$ is an i.i.d. sequence of random variables with law $\mathbb{P}^{(k)}(x,a)$. For a fixed next state $y\in\mathcal{X}$, define
$\xi_{j}^{(y)}:=1\hskip-3.0pt\mathrm{l}_{\{Y_{j}^{(k)}=y\}},$
and, for $n\in\mathbb{N}$, let $S_{n}^{(y)}:=\sum_{j=1}^{n}\xi_{j}^{(y)}$. By the strong law of large numbers, we get
(7.8) $\frac{S_{n}^{(y)}}{n}\rightarrow\mathbb{E}\left[\xi_{j}^{(y)}\right]=\mathbb{P}^{(k)}(x,a)(y)~~\qquad\mathbb{P}_{x_{0},\textbf{a}}-\text{almost\ surely as }n\rightarrow\infty.$
Now, note that
$\mathbb{P}_{t}^{(k)}(x,a)(y)=\frac{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}),X_{s+1}^{(k)})=(x,a,y)\}}}{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}}=\frac{S_{N_{t}(x,a)}^{(y)}}{{N_{t}(x,a)}}$
and the assertion follows with (7.7) and (7.8). ∎
Lemma 7.4.
Let $0<\alpha<1$, and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
(7.9) $\sum^{\infty}_{t=0}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}=\infty,\text{ for all }(x,a)\in\mathcal{X}\times A,\ \mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
Then, we have that
$\|Q_{t}^{*}-Q^{*}\|_{\infty}\leq\frac{\|\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}\qquad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
which implies
${Q}_{t}^{*}(x,a)\rightarrow Q^{*}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A,\qquad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
as $t\rightarrow\infty$.
Proof.
We define $C_{r}:=\max_{(x,a,x^{\prime})\in\mathcal{X}\times A\times\mathcal{X}}|r(x,a,x^{\prime})|<\infty$, and then observe that for all $x\in\mathcal{X}$:
$V(x)=\sup_{\mathbf{a}\in\mathcal{A}}\inf_{\mathbb{P}\in\mathfrak{P}_{x,\textbf{a}}}\mathbb{E}_{\mathbb{P}}\left[\sum^{\infty}_{t=0}\alpha^{t}\cdot r(X_{t},a_{t}(X_{t}),X_{t+1})\right]\leq\sum^{\infty}_{t=0}\alpha^{t}\cdot C_{r}=\frac{C_{r}}{1-\alpha}.$
By Proposition 2.1, for all $(x,a)\in\mathcal{X}\times A$ we then obtain
$Q^{*}(x,a)\leq\max_{a\in A}Q^{*}(x,a)=V(x)\leq\frac{C_{r}}{1-\alpha}$
showing that $\|Q^{*}\|_{\infty}<\infty$. Next, for all $(x,a)\in\mathcal{X}\times A$ we compute
(7.10) $|\mathcal{H}_{t}Q^{*}(x,a)-\mathcal{H}Q^{*}(x,a)|$
$=\left|\min_{\mathbb{P}\in\mathcal{P}_{t}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]-\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]\right|$
$=\bigg|\min_{k=1,\dots,N}\sum_{y\in\mathcal{X}}\left(r(x,a,y)+\alpha\max_{b\in A}Q^{*}(y,b)\right)\mathbb{P}_{t}^{(k)}(x,a)(y)$
$\hskip 170.71652pt-\min_{k=1,\dots,N}\sum_{y\in\mathcal{X}}\left(r(x,a,y)+\alpha\max_{b\in A}Q^{*}(y,b)\right)\mathbb{P}^{(k)}(x,a)(y)\bigg|$
$\leq\max_{k=1,\dots,N}\bigg|\sum_{y\in\mathcal{X}}\left(r(x,a,y)+\alpha\max_{b\in A}Q^{*}(y,b)\right)\left(\mathbb{P}_{t}^{(k)}(x,a)(y)-\mathbb{P}^{(k)}(x,a)(y)\right)\bigg|$
$\leq\left(C_{r}+\alpha\|Q^{*}\|_{\infty}\right)\max_{k=1,\dots,N}\bigg|\sum_{y\in\mathcal{X}}\left(\mathbb{P}_{t}^{(k)}(x,a)(y)-\mathbb{P}^{(k)}(x,a)(y)\right)\bigg|\rightarrow 0\text{ almost surely, as }t\rightarrow\infty$
where the convergence follows by Lemma 7.3 and since $\|Q^{*}\|_{\infty}<\infty$. Next, we get by using Lemma 7.1 and Lemma 7.2 that
$\|Q_{t}^{*}-Q^{*}\|_{\infty}$ $=\|\mathcal{H}_{t}Q_{t}^{*}-\mathcal{H}Q^{*}\|_{\infty}$
$=\|\mathcal{H}_{t}Q_{t}^{*}-\mathcal{H}_{t}Q^{*}+\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
$\leq\alpha\|Q_{t}^{*}-Q^{*}\|_{\infty}+\|\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
implying
$\|Q_{t}^{*}-Q^{*}\|_{\infty}\leq\frac{\|\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}$
which by (7.10) converges against 0 almost surely as $t\rightarrow\infty$. ∎
The following lemma is a result from stochastic approximation theory compare also [49, Lemma 1], [16] [23, Theorem 1], [51, Lemma 12], and [57, Lemma 3] for further reference.
@@
    There exists $\delta\in(0,1)$ such that $\|\mathbb{E}_{\mathbb{P}_{0}}\left[F_{t}(\cdot,\ \cdot)~\middle|~\mathcal{G}_{t}\right]\|_{\infty}\leq\delta\|\Delta_{t}\|_{\infty}+c_{t}$  $\mathbb{P}_{0}$-almost surely for all $t\in\mathbb{N}_{0}$, where $c_{t}$ converges to zero  $\mathbb{P}_{0}$-almost surely.
@@
Let $Z$ be a random variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ satisfying $m\leq Z\leq M$ for some $\infty<m\leq M<\infty$. Then, we have
@@
To keep the paper self-contained, we further state the so called Azuma–Hoeffding Inequality, see, e.g., [33] or [34].
Lemma 7.7 ([5]).
Let $(M_{t},\mathcal{F}_{t})_{t\in\mathbb{N}}$ be a martingale with respect to the filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. Assume that the martingale differences are almost surely bounded, i.e., there exist constants $c_{t}>0$ such that
$|M_{t}-M_{t-1}|\leq c_{t}\quad\text{a.s. for all }t\in\mathbb{N}$
Then for any $\varepsilon>0$, $n\in\mathbb{N}$ we have
$\mathbb{P}\!\left(|M_{n}-M_{0}|\geq\varepsilon\right)\leq 2\exp\!\left(-\frac{\varepsilon^{2}}{2\sum_{t=1}^{n}c_{t}^{2}}\right).$
Moreover, we make use of the following concentration inequality for empirical distributions.
Lemma 7.8 ([63], Theorem 2.1).
Let $\mathbb{Q}$ be a probability distribution on $\mathcal{X}$. Let $\widehat{\mathbb{\mathbb{Q}}}_{n}$ denote the empirical distribution of $n$ samples from $\mathbb{Q}$. Then for all $\varepsilon>0$
$\mathbb{P}\left(\|\widehat{\mathbb{\mathbb{Q}}}_{n}-\mathbb{Q}\|_{1}\geq\varepsilon\right)\leq\left(2^{|\mathcal{X}|}-2\right)\exp\left(-\frac{n\varepsilon^{2}}{2}\right).$
@@ 7.2. Proofs of the results from Section 2, Section 3, and Section 4
Note that the equality $\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)$ directly follows by definition of the operator $\mathcal{T}$, and it only remains to show $\mathcal{T}V=V$. To this end, we aim at applying [38, Theorem 3.1] with $p=0$ and verify its assumptions, i.e., we verify [38, Assumption 2.2] and [38, Assumption 2.4] by showing the following four properties.
@@
    $(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a)=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\}$
@@
To see that (i) holds, we note that $\mathcal{P}$ is by definition non-empty. It is also compact for all $(x,a)\in\mathcal{X}\times A$ since the set $\mathcal{P}(x,a)$ is finite. To show the upper hemicontinuity, we consider some $(x,a)\in\mathcal{X}\times A$ and we let $(x_{n},a_{n},\mathbb{P}_{n})_{n\in\mathbb{N}}\subseteq\operatorname{Gr}(\mathcal{P}):=\{(x,a,\mathbb{P})~|~(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)\}$ be a sequence such that $(x_{n},a_{n})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$. Since both $\mathcal{X}$ and $A$ are finite, we then have $(x_{n},a_{n})=(x,a)$ for $n$ large enough. Hence, $\mathbb{P}_{n}\in\mathcal{P}(x,a)$ for $n$ large enough. As $\mathcal{P}(x,a)$ is finite we also have that $\mathbb{P}_{n}=\mathbb{P}^{(i)}(x,a)$ for infinitely many indices $n$ for some $i\in\{1,\dots,N\}$. Hence, we can find a subsequence $(\mathbb{P}_{n_{k}})_{k\in\mathbb{N}}$ such that $\mathbb{P}_{n_{k}}\rightarrow\mathbb{P}^{(i)}(x,a)\in\mathcal{P}(x,a)$ weakly as $k\rightarrow\infty$ and the upper hemicontinuity of $\mathcal{P}$ follows by the characterization provided in [2, Theorem 17.20].
@@
and we obtain, since $(x_{n},a_{n})=(x,a)$ for $n$ large enough that
$\mathbb{P}_{n}=\mathbb{P}^{(i_{\mathbb{P}})}(x_{n},a_{n})=\mathbb{P}^{(i_{\mathbb{P}})}(x,a)=\mathbb{P}$
for $n$ large enough that and hence $\mathbb{P}_{n}\rightarrow\mathbb{P}$ weakly which shows the lower hemicontinuity of $\mathcal{P}$ according to [2, Theorem 17.21].
@@
Finally, (iv) follows since $C_{P}=1$ and as we have $0<\alpha<1$ by assumption. ∎
@@
To establish convergence of the $Q$-learning algorithm from Algorithm 1, we apply Lemma 7.5, and to verify the assumptions of Lemma 7.5 we first note that for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ the update rule from Algorithm 1, expressed in (3.3), can be written equivalently as
@@
(7.13) $\gamma_{t}(x,a,X_{t}):=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{N_{t}(x,a)}.$
Next, for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ we define
$\Delta_{t}(x,a):$ $=Q_{t}(x,a)-Q_{t}^{*}(x,a),$
$F_{t}(x,a):$ $=\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right)1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}.$
@@
$k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,...,N}\mathbb{E}_{\mathbb{P}_{t}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right].$
Once the three assumptions of Lemma 7.5 are verified, it follows that $\Delta_{t}(x,a)\rightarrow 0$ $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely as $t\rightarrow\infty$ and hence with Lemma 7.4
$\|Q_{t}-Q^{*}\|_{\infty}\leq\|Q_{t}-Q_{t}^{*}\|_{\infty}+\|Q_{t}^{*}-Q^{*}\|_{\infty}=\|\Delta_{t}\|_{\infty}+\|Q_{t}^{*}-Q^{*}\|_{\infty}\rightarrow\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$
and thus
$Q_{t}(x,a)\rightarrow Q^{*}(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$
@@
$\Delta_{t+1}(x,a)=$ $Q_{t}(x,a)-Q_{t}^{*}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)\Big)$
@@
$+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)+Q_{t}^{*}(x,a)-Q_{t}^{*}(x,a)\Big)$
@@
$+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\Big)$
@@
where we used in the last step that the indicator $1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}$ appears both in the definition of $F_{t}$ and $\gamma_{t}$ and that the square of the indicator function is the indicator function itself. Consider the filtration $(\mathcal{G}_{t})_{t\in\mathbb{N}}$ with
$\mathcal{G}_{t}:=\sigma(X_{1}^{(1)},X_{2}^{(1)},...,X_{t}^{(N)}),\ t\in\mathbb{N},$
and $\mathcal{G}_{0}:=\{\varnothing,\Omega^{N}\}$ being the trivial sigma-algebra. Then, by definition we have for all $t\in\mathbb{N}$ and for all $(x,a)\in\mathcal{X}\times A$ that the random variables $\Delta_{t}(x,a),\ \gamma_{t}(x,a),\ F_{t-1}(x,a)$ are $\mathcal{G}_{t}$-measurable.
We now define the random operator
(7.14) $\mathcal{X}\times A\ni(x,a)\mapsto\widetilde{\mathcal{H}}_{t}q(x,a):=\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right],$
where $k_{t}^{*}$ is the index selected by the empirical minimization in Algorithm 1; expectations are taken under the corresponding true kernel $\mathbb{P}^{(k_{t}^{*})}$ to quantify the kernel estimation error between $\mathbb{P}_{t}^{(k_{t}^{*})}$ and $\mathbb{P}^{(k_{t}^{*})}$. Further, we define $C_{r}:=\max_{y_{0},y_{1}\in\mathcal{X},\ b\in A}|r(y_{0},b,y_{1})|<\infty$ and then, analogous to the proof of Lemma 7.4, we obtain $\|Q_{t}\|_{\infty}<\frac{C_{r}}{1-\alpha}<\infty$. Next, let $(x,a)\in\mathcal{X}\times A,$ and $t\in\mathbb{N}$, and note that
$\left|\widetilde{\mathcal{H}}_{t}Q_{t}(x,a)-{\mathcal{H}}_{t}Q_{t}(x,a)\right|$
$=\left|\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]-\mathbb{E}_{\mathbb{P}_{t}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]\right|$
$\leq\max_{k=1,\dots,N}\left|\mathbb{E}_{\mathbb{P}^{(k)}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]-\mathbb{E}_{\mathbb{P}_{t}^{(k)}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]\right|$
$\leq\left(C_{r}+\alpha\|Q_{t}\|_{\infty}\right)\max_{(x,a)\in\mathcal{X}\times A}\max_{k=1,\dots,N}\sum_{y\in\mathcal{X}}\left|\mathbb{P}_{t}^{(k)}(x,a)(y)-\mathbb{P}^{(k)}(x,a)(y)\right|=:c_{t}$
with $c_{t}\rightarrow 0$ $\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$ by Lemma 7.3 and since $\|Q_{t}\|_{\infty}<\infty$. Then, we have by the definition of $\mathbb{P}_{x_{0},\textbf{a}}$ in (3.4) $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely that